If f:R^2 to R defined by f(x,y)={((x^2 y^3)/(x^4 + y^2)) if (x,y) # (0,0)
0 if (x,y) = (0,0) . Find the directional derivative of f(x,y) at (0,0)?
1
Expert's answer
2020-05-18T19:44:42-0400

for (x,y)"\neq"(0,0),


f(x,y) = "\frac{x^2 y^3}{x^4+y^2}"


the directional derivative of f at A=(a,b) in the direction of U=(u1,u2) is


D = "\lim_{t \to o} \frac{f(A+tU)-f(A)}{t} = \lim_{t \to o} \frac{f(tu_1 , tu_2)-f(0,0)}{t}"


"\lim_{t \to o} \frac{t^5 u_1^2 u_2^3}{t^3 (t^2 u_1 ^4+ u_2 ^2)}"


"\lim_{t \to o} \frac{t^2 u_1 ^2 u_2 ^3}{t^2 u_1 ^4+ u_2 ^2}" = 0